OpenStudy (anonymous):

Find the center, radius, and intercepts. 2x^2+2y^2+8x+7=0

OpenStudy (anonymous):

I need help on putting this in Standard Form. From there, I believe I can do the rest.

OpenStudy (anonymous):

you gotta complete the square twice once for the x terms and once for the y terms

OpenStudy (anonymous):

I did and got 2(x+2)^2+2(y)^2=1 I'm not sure if this is correct though.

OpenStudy (anonymous):

\[2x^2+2y^2+8x+7=0\] subtract \(7\)\[2x^2+2y^2+8x=-7\] divide by 2 to make the leading coefficient 1 and start with \[x^2+4x+y^2=-\frac{7}{2}\]

OpenStudy (anonymous):

it is not correct, you can't complete the square without making the leading coefficient one

OpenStudy (anonymous):

standard form \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

Oh! Thank you!

OpenStudy (anonymous):

you good from there? how to complete the square i mean?

OpenStudy (anonymous):

Yes! I just have one last question, how can you tell if you have x-int and y-int?

OpenStudy (anonymous):

x intercept if what you get when you put \(y=0\)

OpenStudy (anonymous):

you can do that without putting it in standard form, just get rid of the y and solve for \(x\)

OpenStudy (anonymous):

but you need the center and radius anyway, so you have no choice but to complete the square

OpenStudy (anonymous):

Oh yeah! :D Thank you!